OpenStudy (anonymous):

Please Help, I need help with word problem. Write the equation for this situation. The cost, T, in hundreds of dollars, at one community college is 5 more than 1/2 the number of credits, c.

OpenStudy (anonymous):

OK, there a few tricks for converting the phrase to a mathematical equation. 1. 'is' is equivalent to '='

OpenStudy (anonymous):

I can solve it if it is out of word problem i always have a hard time i'm spending all my time reading it over and over

OpenStudy (anonymous):

We will break convert the phrase: 'The cost, T, in hundreds of dollars, at one community college is 5 more than 1/2 the number of credits, c.' Forget the modifying phrase 'in hundreds of dollars' and where the cost is: 'at one community college' Now, you have: 'The cost, T, is 5 more than 1/2 the number of credits, c.

OpenStudy (anonymous):

T is 5 more than 1/2 the number of credits c

OpenStudy (anonymous):

so would it be T=5+1/2c

OpenStudy (anonymous):

T = (5 more than 1/2)c

OpenStudy (anonymous):

Well, you could be correct. There seems to be a bit ambiguity, at least to me

OpenStudy (anonymous):

So, you have said 'T = 5 + 1/2*c' and perhaps 'T = (5+1/2) * c' , But now that I think about, your answer seems best.

OpenStudy (anonymous):

nice work

OpenStudy (anonymous):

thank you so much i am 40yr old and finally understand word problems

OpenStudy (anonymous):

That's awesome! It seems that you've got the hang of it ;) Surely, learning does just take time.